OpenStudy (anonymous):

3 ln X + ln 24= ln 3

OpenStudy (superdavesuper):

Use these two: ln(a) + ln(b) = ln (a*b) and a*ln(b) = ln(b)^a

OpenStudy (isaiah.feynman):